How does the equation for valuing a bond change if semiannual payments are made? Find the value of a 10-year, semiannual payment, 10 percent coupon bond if nominal rd =13
1
Expert's answer
2020-10-13T11:08:35-0400

"Bond Price= C \times \frac {1- (1+r)-n} {r} + \frac {F} {(1+r)n}"


When semi annual payments are made, n which is the number of periods to maturity is doubled while r which is the yield of the bond is halved


"n= 10 \times 2= 20"


"FV= 1000"

"PMT= \frac {100} {2}= 50"


"r= \frac {13} {2}= 6.5%"


Therefore PV = $ 834.72